1) Suppose that |a|≤ 2 and |b|≤3.
a) What is the largest possible value of |a+b|?
b) What is the largest possible value of |a+b| if a and b have opposite signs?

Thank you for your help

Answer by stanbon(75887)

1) Suppose that |a|≤ 2 and |b|≤3.
-2<= a <=2
-3<= b <=3
------------
a) What is the largest possible value of |a+b|?
Ans: |a+b|= |2+3| = 5
-----
b) What is the largest possible value of |a+b| if a and b have opposite signs?
Ans: b = 3; a is negative but infinitesimly small
|a+b| has limit 3
------------------
Cheers,
Stan H.